OpenStudy (anonymous):

Find the slope intercept equation for the line with the indicated slope and y intercept. Slope -4, y-intercept (0, -3/5)

OpenStudy (amorfide):

 m= slope y1=y coordinate x1= x coordinate

OpenStudy (anonymous):

y-y1=m(x-x1) y--(3/5) = -4(x-0) y+(3/5)=-4x y=-4x-(3/5)